# Weak convergence of Bayes estimators under general loss functions

### v1

@@ 6.2 Multinomial–Dirichlet model
As in Section 5.3, consider the multinomial model $\{P_{p}=\mathrm{Mult}(1,p):p\in[0,1]^{d},\ \left\lVert p\right\rVert_{1}=1\}$ on $\mathcal{X}=\{y_{1},\dots,y_{d}\}$ with $y_{j}$ the $j$th unit vector in $\mathbb{R}^{d}$. Now we employ a Dirichlet prior $\mathrm{Dir}(\alpha)$, and then the posterior is $\mathrm{Dir}(\alpha+\sum_{i=1}^{n}x_{i})$, given ${\boldsymbol{X}}=(x_{1},\ldots,x_{n})$.
The squared 2-Wasserstein distance between two multinomial measures $P_{p}$ and $P_{\vartheta}$ can be expressed explicitly in terms of their coordinates (Lemma 2). In particular, $W_{2}^{2}(P_{p},P_{\vartheta})\ =\ 1-\sum_{i=1}^{d}\min(p_{i},\vartheta_{i}).$ Since $\{p\in[0,1]^{d}:\left\lVert p\right\rVert_{1}=1\}$ has empty interior in $\mathbb{R}^{d}$, we reparametrize as $\Theta=\left\{(p_{1},\dots,p_{d-1})\in[0,1]^{d-1}:\sum_{i=1}^{d-1}p_{i}\leq 1\right\},$ with $\tilde{P}_{a}=P_{(a,\,1-\left\lVert a\right\rVert_{1})}$ for $a\in\Theta$. The loss becomes $\ell(a,b)=W_{2}^{2}(\tilde{P}_{a},\tilde{P}_{b})=\sum_{i=1}^{d-1}a_{i}+b_{i}-\min(a_{i},b_{i})=\left\lVert a-b\right\rVert_{1},\;a,b\in\Theta,$ i.e., the 2-Wasserstein distance coincides with the total variation distance. The inverse of the Fisher information matrix at $\vartheta_{0}=(p_{1},\dots,p_{d-1})\in\Theta^{\circ}$ is given by
@@
The limiting process is then $Z(t)=\int_{\mathbb{R}^{d-1}}\left\lVert t-h\right\rVert_{1}\,\mathop{}\!\mathrm{d}\mathcal{N}_{d-1}\!\left(Y,I^{-1}_{\vartheta_{0}}\right)(h),$ which is minimized at $t=Y$ and the Bayes estimator is the vector of component wise marginal posterior medians. If ${\boldsymbol{X}}=x$, the $k$th marginal of the posterior $\mathrm{Dir}(\alpha+\sum_{i=1}^{n}x_{i})$ is $\mathrm{Beta}(\tilde{\alpha}_{k},\sum_{j\neq k}\tilde{\alpha}_{j})$, with $\tilde{\alpha}_{k}=\alpha_{k}+\sum_{i=1}^{n}x_{i,k}$, so the $k$th component of $\hat{\theta}_{n}$ is the median of this Beta distribution.
@@
In the simulation, we set $d=3$, $\alpha=(1,1,1)$, and $\vartheta_{0}=(1/3,1/3)$. Figure 3 shows the 2-Wasserstein distance between the empirical distribution of $n^{1/2}(\hat{\theta}_{n}-\vartheta_{0})$ (over $R=100$ Monte Carlo repetitions) and $\mathcal{N}_{d-1}(0,I_{\vartheta_{0}}^{-1})$ as a function of $n$. The Wasserstein distance was obtained by $M=2000$ samples of the Bayes estimator and gaussian random variables. One can see that the distance converges to zero. As the 2-Wasserstein distance metrizes weak converge this yields that the Bayes estimator converges to the normal limit $\mathcal{N}_{d-1}\!\left(0,I^{-1}_{\vartheta_{0}}\right)$.
@@
Figure 3: 2-Wasserstein distance between the empirical distribution of $n^{1/2}(\hat{\theta}_{n}-\vartheta_{0})$ and $\mathcal{N}_{d-1}(0,I_{\vartheta_{0}}^{-1})$ for $d=3$, $\alpha=(1,1,1)$ and $\vartheta_{0}=(1/3,1/3)$. Each point is based on $R=100$ Monte Carlo repetitions of calculating the Wasserstein distance based on $M=2000$ samples.
@@ Appendix A Various types of differentiability
In this section we give an overview over different concepts of differentiability considered in this paper and put them into perspective.
@@ A.1 Hadamard derivative
Let $\phi:D\to E$ be a function between two Banach spaces $D$ and $E$. We say $\phi$ is $p$-th order Hadamard directionally differentiable at $\vartheta_{0}\in D^{\circ}$ in direction $h\in D$, if there exists a map $\phi^{\prime}_{\vartheta_{0}}:D\to E$ such that
@@
for any sequences $r_{n}\to 0,r_{n}>0$ and $h_{n}\to h$ and for $p>0$. Here the derivative $\phi^{\prime}_{\vartheta_{0}}(\cdot)$ does not have to be linear, but is by definition continuous at $h$. If $p=1$ and $\phi^{\prime}_{\vartheta_{0}}(\cdot)$ is linear we say that $\phi$ is Hadamard differentiable at $\vartheta_{0}\in D^{\circ}$ in direction $h\in D$. In Averbukh and Smolyanov, (1968, Theorem 3.2) it is shown that for any compact set $K\subset D$, the Hadamard differentiability at $\vartheta_{0}\in D^{\circ}$ in direction $h$ for all $h\in K$ is equivalent to the compactly differentiability at $\vartheta_{0}$, i.e. there exists a continuous, linear map $\phi^{\prime}_{\vartheta_{0}}:D\to E$ with
@@
Similarly, we say $\phi$ is $p$-th order compactly directionally differentiable at $\vartheta_{0}$, if there exists a map $\phi^{\prime}_{\vartheta_{0}}:D\to E$ such that
@@
Proposition 4.
(Shapiro,, 1990, Proposition 3.3). Let $p>0$. The $p$-th order Hadamard directional differentiability of $\phi$ at $\vartheta_{0}$ in direction $h$ for all $h\in D$ implies $p$-th order compact directional differentiability of $\phi$ at $\vartheta_{0}$. Conversely, if $\phi$ is $p$-th order compactly directionally differentiability with derivative $\phi^{\prime}_{\vartheta_{0}}$, and if $\phi^{\prime}_{\vartheta_{0}}$ is sequentially continuous, then $\phi$ is $p$-th order Hadamard directionally differentiable at $\vartheta_{0}$.
@@
Another concept, which is stronger than Hadamard differentiability is Fréchet differentiability. We say $\phi$ is Fréchet differentiable at $\vartheta_{0}\in D^{\circ}$, if if there exists a continuous and linear map $\phi^{\prime}_{\vartheta_{0}}:D\to E$ such that
@@
Connection to Condition (iii) from Proposition 3. In our setting we consider $D=\Theta\times\Theta$ for some Banach space $\Theta$ and are only interested in differentiability at the point $a_{0}=(\vartheta_{0},\vartheta_{0})$ for $\vartheta_{0}\in\Theta^{\circ}$. Consider a loss function $\ell:\Theta\times\Theta\to[0,\infty)$ and set $\phi(a)=\ell(a_{1},a_{2})$ for $a=(a_{1},a_{2})\in\Theta\times\Theta$. Condition (iii) from Proposition 3 requires that for $\Theta\subseteq\mathbb{R}^{d}$ that there exists $\phi^{\prime}_{\vartheta_{0}}:D\to[0,\infty)$ such that for all $h\in\mathbb{R}^{d}$ and compact set $K\subseteq\mathbb{R}^{d}$
@@
$\zeta(t,h)>0,\text{if }t\neq h,\quad\zeta(\lambda t,\lambda h)=\lambda^{p}\zeta(t,h)\quad\text{and}\quad\zeta(t+a,h+a)=\zeta(t,h),\quad\forall a,t,h\in\mathbb{R}^{d},\lambda>0$
@@
For $d=2$ and $\vartheta_{0}=(0,0)$ the map $t\mapsto Z(t)$ is visualized for fixed $Y=(1,1)$. Here $Z$ measures the expected curvature of $\ell$ at $\vartheta_{0}$.
Figure 4: Plot of $(t_{1},t_{2})\mapsto Z(t_{1},t_{2})$ for $Y=(1,1)$. The green arrow visualizes the vector $(1,1)$.
A.2 $L^{p}$ derivative
@@
Let $\mathcal{M}=\{f_{\vartheta}\in L^{p}:\vartheta\in\Theta\}$, with $\Theta\subseteq\mathbb{R}^{d}$, be a parametric family of functions. We say the family $\mathcal{M}$ is $L^{p}$ differentiable at $\vartheta_{0}\in\Theta^{\circ}$, if there exists $\dot{f}_{\vartheta_{0}}:\mathcal{X}\to\mathbb{R}^{d}$ with $\left\lVert\dot{f}_{\vartheta_{0}}\right\rVert\in L^{p}$ and
@@
as $\vartheta\to\vartheta_{0}$. In this form we can see directly that this is the Fréchet differentiability in the space $L^{p}$. We consider three examples of $L^{p}$ differentiability: $L^{2}$ differentiability of the square root of a density, $L^{1}$ differentiability of a density and $L^{1}$ differentiability of a distribution function.
Differentiability in quadratic mean. We say that a parametric family of probability distributions $\{P_{\vartheta}:\vartheta\in\Theta\}$, dominated by some measure $\mu$ on $\mathcal{X}$, is differentiable in quadratic mean at $\vartheta_{0}\in\Theta^{\circ}$, if there exists a vector valued function $\ell^{\prime}_{\vartheta_{0}}:\mathcal{X}\to\mathbb{R^{d}}$ such that
@@
Note that the differentiability in quadratic mean is the same concept as the $L^{2}$ differentiability of $\{p^{1/2}_{\vartheta}:\vartheta\in\Theta\}$. Differentiability in quadratic mean implies the existence of the Fisher matrix and
@@
$L^{1}$ differentiability. Differentiability in quadratic mean at $\vartheta_{0}$ implies that the $L^{1}$ differentiability of $p_{\vartheta}$ at $\vartheta_{0}$ and the derivative is given by $\dot{p}_{\vartheta_{0}}=\ell^{\prime}_{\vartheta_{0}}p_{\vartheta_{0}}$, see Liese and Miescke, (2008, Proposition 1.110). So the differentiability in quadratic mean is a stronger property than $L^{1}$ differentiability of the density.
The following proposition gives conditions for $L^{1}$ differentiability of the distribution function $F_{\vartheta}(t)=\int_{-\infty}^{t}p_{\vartheta}(x)dx$. Under the first condition the result was shown in Proposition B.2 from the supplement of Bernton et al., 2019b . We extend this by allowing $\mathcal{X}$ to be unbounded.
@@
    For any $x\in\mathcal{X}$ the map $\vartheta\mapsto p_{\vartheta}(x)$ is continuously differentiable at $\vartheta_{0}\in\Theta^{\circ}$ with derivative $\dot{p}_{\vartheta_{0}}$.
@@
By differentiability in quadratic mean it follows from Liese and Miescke, (2008, Proposition 1.110) that the $L^{1}$ derivative of $p_{\vartheta}$ at $\vartheta_{0}$ is given by $\dot{p}_{\vartheta_{0}}=\ell^{\prime}_{\vartheta_{0}}p_{\vartheta_{0}}$. Define
$\mathcal{F}(t)=\int_{-\infty}^{t}\left\lvert p_{\vartheta}(x)-p_{\vartheta_{0}}(x)-\left\langle\vartheta-\vartheta_{0},\,\dot{p}_{\vartheta_{0}}(x)\right\rangle\right\rvert\mathop{}\!\mathrm{d}\mu(x)=\left\lVert\left(p_{\vartheta}-p_{\vartheta_{0}}-\left\langle\vartheta-\vartheta_{0},\,\dot{p}_{\vartheta_{0}}\right\rangle\right)\mathbb{I}(\cdot\leq t)\right\rVert_{L^{1}}.$
@@
$\left\lVert\vartheta-\vartheta_{0}\right\rVert^{-1}\int_{\mathcal{X}}\left\lvert F_{\vartheta}(t)-F_{\vartheta_{0}}(t)-\left\langle\vartheta-\vartheta_{0},\,\dot{F}_{\vartheta_{0}}(t)\right\rangle\right\rvert\mathop{}\!\mathrm{d}\mu(t)\leq\int_{\mathcal{X}}\frac{\mathcal{F}(t)}{\left\lVert\vartheta-\vartheta_{0}\right\rVert}\mathop{}\!\mathrm{d}\mu(t).$
@@
$0\leq\lim_{\vartheta\to\vartheta_{0}}\frac{\mathcal{F}(t)}{\left\lVert\vartheta-\vartheta_{0}\right\rVert}\leq\lim_{\vartheta\to\vartheta_{0}}\frac{\left\lVert p_{\vartheta}-p_{\vartheta_{0}}-\left\langle\vartheta-\vartheta_{0},\,\dot{p}_{\vartheta_{0}}\right\rangle\right\rVert_{L^{1}}}{\left\lVert\vartheta-\vartheta_{0}\right\rVert}=0$
by $L^{1}$ differentiability of $p_{\vartheta}$ at $\vartheta_{0}$. If $\mathcal{X}$ is bounded, we are done. If condition (ii) holds, we use the dominated convergence theorem with the integrable upper bound
$\frac{\mathcal{F}(t)}{\left\lVert\vartheta-\vartheta_{0}\right\rVert}\leq\sup_{\tilde{\vartheta}\in B_{\delta}(\vartheta_{0})}\left\lVert\left\lVert\dot{p}_{\tilde{\vartheta}}-\dot{p}_{\vartheta_{0}}\right\rVert\right\rVert_{L^{1}}$
@@
As in Section 5.2 this is closely related to the differentiability of 1-Wasserstein distance.

### v2

@@ 6.2 Multinomial–Dirichlet model
Consider the multinomial model $\{P_{p}=\mathrm{Mult}(1,p):p\in[0,1]^{d},\ \left\lVert p\right\rVert_{1}=1\}$ on $\mathcal{X}=\{y_{1},\dots,y_{d}\}$ with $y_{j}$ the $j$th unit vector in $\mathbb{R}^{d}$, as in Section 5.3. Now we employ a Dirichlet prior $\mathrm{Dir}(\alpha)$, and then the posterior is $\mathrm{Dir}(\alpha+\sum_{i=1}^{n}x_{i})$, given ${\boldsymbol{X}}=(x_{1},\ldots,x_{n})$.
The squared 2-Wasserstein distance between two multinomial measures $P_{p}$ and $P_{\vartheta}$ can be expressed in terms of their coordinates (Lemma 2 in Supplement C.6). In particular, $W_{2}^{2}(P_{p},P_{\vartheta})\ =\ 1-\sum_{i=1}^{d}\min(p_{i},\vartheta_{i}).$ Since $\{p\in[0,1]^{d}:\left\lVert p\right\rVert_{1}=1\}$ has empty interior in $\mathbb{R}^{d}$, we reparametrize as $\Theta=\left\{(p_{1},\dots,p_{d-1})\in[0,1]^{d-1}:\sum_{i=1}^{d-1}p_{i}\leq 1\right\},$ with $\tilde{P}_{a}=P_{(a,\,1-\left\lVert a\right\rVert_{1})}$ for $a\in\Theta$. The loss becomes
$\ell(a,b)=W_{2}^{2}(\tilde{P}_{a},\tilde{P}_{b})=\sum_{i=1}^{d-1}\bigl(a_{i}+b_{i}-\min\{a_{i},b_{i}\}\bigr)=\left\lVert a-b\right\rVert_{1},\;a,b\in\Theta,$
i.e., the 2-Wasserstein distance coincides with the total variation distance. The inverse of the Fisher information matrix at $\vartheta_{0}=(p_{1},\dots,p_{d-1})\in\Theta^{\circ}$ is given by
@@
The limiting process is then $Z(t)=\int_{\mathbb{R}^{d-1}}\left\lVert t-h\right\rVert_{1}\,\mathop{}\!\mathrm{d}\mathcal{N}_{d-1}\!\left(Y,I^{-1}_{\vartheta_{0}}\right)(h),$ which is minimized at $t=Y$ and the Bayes estimator is the vector of component wise marginal posterior medians. If ${\boldsymbol{X}}=x$, the $k$-th marginal of the posterior $\mathrm{Dir}(\alpha+\sum_{i=1}^{n}x_{i})$ is $\mathrm{Beta}(\tilde{\alpha}_{k},\sum_{j\neq k}\tilde{\alpha}_{j})$, with $\tilde{\alpha}_{k}=\alpha_{k}+\sum_{i=1}^{n}x_{i,k}$, so the $k$-th component of $\hat{\theta}_{n}$ is the median of this Beta distribution.
@@
In the simulation, we set $d=3$, $\alpha=(1,1,1)$, and $\vartheta_{0}=(1/3,1/3)$. Figure 3 shows the 2-Wasserstein distance between the empirical distribution of $n^{1/2}(\hat{\theta}_{n}-\vartheta_{0})$ and $\mathcal{N}_{d-1}(0,I_{\vartheta_{0}}^{-1})$ as a function of $n$. For each $n$, the 2-Wasserstein distance between the empirical distribution of the scaled Bayes estimator and the limiting normal distribution was computed using the function wasserstein in the R package transport Schuhmacher et al., (2024). The computation was based on $10^{4}$ drawn from each distribution and averaged over $M=100$ Monte Carlo repetitions, following the approach of Sommerfeld et al., (2019). One can see that the distance rapidly converges to zero. Since the 2-Wasserstein distance metrizes weak converge, this implies that the Bayes estimator, after proper scaling, converges in distribution to the normal limit $\mathcal{N}_{d-1}\!\left(0,I^{-1}_{\vartheta_{0}}\right)$.
@@
Figure 3: 2-Wasserstein distance between the empirical distribution of $n^{1/2}(\hat{\theta}_{n}-\vartheta_{0})$ and $\mathcal{N}_{d-1}(0,I_{\vartheta_{0}}^{-1})$ for $n\in\{10,25,50,100,225,500,1000\}$. Here $\hat{\theta}_{n}$ is the Bayes estimator, $d=3$, $\alpha=(1,1,1)$ and $\vartheta_{0}=(1/3,1/3)$. Each point represents the average over $M=100$ Monte Carlo repetitions, where the Wasserstein distance is computed from $10^{4}$ samples from each distribution. Both axes are shown on a logarithmic scale.
7 Discussion
In this paper, we study the weak convergence of the rescaled Bayes estimator $\delta_{n}^{-1}(\hat{\theta}_{n}-\vartheta_{0}),$ with $\hat{\theta}_{n}$ defined through a general loss function $\ell(\cdot,\cdot)$, under the frequentist setting where observations $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}P_{\vartheta_{0}},$ for a fixed parameter $\vartheta_{0}\in\Theta$, a subset of a metric vector space. The scaling $\delta_{n}$ is determined by the posterior contraction rate; in the regular parametric models considered here, $\delta_{n}=1/\sqrt{n}$. The proof strategy, rooted in the classical works of LeCam, (1953); Chao, (1970); Ibragimov and Has minskiĭ, (1973), consists of two main steps: (a) establishing weak convergence of the posterior risk as a stochastic process indexed by the local parameter $t=\delta_{n}^{-1}(\vartheta-\vartheta_{0})$, with $\vartheta$ ranging over a neighborhood of $\vartheta_{0}$; and (b) applying the argmax continuous mapping theorem (see e.g. (van der Vaart,, 1998, Theorem 5.56) and (van der Vaart and Wellner,, 2023, Theorem 3.2.2)).
This framework is potentially applicable beyond the setting studied here and could provide a route to analyzing Bayes estimators in more general models. In particular, step (a), namely the convergence of the posterior risk, depends crucially on an approximation to the posterior distribution, such as BvM-type results, or Laplace approximations. Such approximations have been established in a variety of settings, including non-regular models (Bochkina and Green,, 2014; Ghosal and Samanta,, 1997), misspecified models (Kleijn and van der Vaart,, 2012), semiparametric models (Bickel and Kleijn,, 2012; Castillo,, 2012; Castillo and Rousseau,, 2015), nonparametric models (Castillo and Nickl,, 2014; Ghosal and van der Vaart,, 2017), and high-dimensional models (Katsevich,, 2025; Spokoiny,, 2025), see also (Durante et al.,, 2024) for non-Gaussian approximations. To derive the convergence of posterior risk in step (a) from the posterior approximations, and to establish uniform tightness (or an appropriate alternative) of $\delta_{n}^{-1}(\hat{\theta}_{n}-\vartheta_{0})$, one additionally needs a strengthened form of posterior contraction, as in Proposition 7 (in Supplement B). Thus, a systematic treatment of these extensions would require substantial additional work and lies beyond the scope of the present paper. We nevertheless briefly discuss how our analysis may extend to misspecified models.
Following Kleijn and van der Vaart, (2012), suppose now that $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}P_{0},$ where $P_{0}\notin\{P_{\vartheta}:\vartheta\in\Theta\subseteq\mathbb{R}^{d}\}.$ Define the pseudo-true parameter
$\vartheta_{*}=\arg\min_{\vartheta\in\Theta}\mathrm{KL}(P_{0}\mid P_{\vartheta}),$
where $\mathrm{KL}$ denotes the Kullback–Leibler divergence, and assume that this minimizer is unique. We further assume that the matrix
$V_{\vartheta_{*}}=\nabla_{\vartheta}^{2}\mathrm{KL}(P_{0}\mid P_{\vartheta})\big|_{\vartheta=\vartheta_{*}}$
is nonsingular. Under these conditions, Theorem 2.1 of Kleijn and van der Vaart, (2012) yields a BvM-type result: the posterior distribution of $\sqrt{n}(\hat{\theta}_{n}-\vartheta_{*})$ converges in total variation distance to a normal distribution with covariance matrix $V_{\vartheta_{*}}^{-1}$.
Moreover, we consider the loss function that is bounded by a polynomial, a stricter requirement than the exponential growth in Definition 1. This may be viewed as a price of model misspecification. For such losses, we can extend Proposition 7 to this misspecified setup. As a consequence, we may establish, much as in Proposition 3, the convergence of the posterior risk in step (a) to a nondegenerate limit process $Z(t)$ in $\ell^{\infty}(K)$ for every compact set $K\subseteq\mathbb{R}^{d}$, as well as the uniform tightness of $\sqrt{n}(\hat{\theta}_{n}-\vartheta_{*})$, which is a key ingredient for step (b). The limit process of posterior risks would take the form
$Z(t)=\int_{\mathbb{R}^{d}}\ell_{0}(t,h)\,\mathop{}\!\mathrm{d}\mathcal{N}_{d}(Y,V^{-1}_{\vartheta_{*}})(h),$
where $\ell_{0}$ is a directional Hadamard derivative of the loss $\ell$, $Y\sim\mathcal{N}_{d}(0,J_{\vartheta_{*}})$ with
$J_{\vartheta_{*}}=V_{\vartheta_{*}}^{-1}\left(\int_{\mathcal{X}}\ell_{\vartheta_{*}}^{\prime}(x){\ell_{\vartheta_{*}}^{\prime}(x)}^{\intercal}\,\mathop{}\!\mathrm{d}P_{0}(x)\right)V_{\vartheta_{*}}^{-1},$
and recall that $\ell_{\vartheta_{*}}^{\prime}$ denotes the score function. This suggests that an extension of Theorem 2 to misspecified models should be feasible. We leave a careful investigation of this extension, as well as of the other directions mentioned above, to future work.
@@ Appendix A Various types of differentiability
In this supplement, we give an overview over different concepts of differentiability considered in this paper, and put them into perspective.
@@ A.1 Hadamard derivative
Let $\phi:D\to E$ be a function between two Banach spaces $D$ and $E$. We say $\phi$ is $p$-th order Hadamard directionally differentiable at $\vartheta_{0}\in D^{\circ}$ in direction $h\in D$, if there exists a mapping $\phi^{\prime}_{\vartheta_{0}}:D\to E$ such that
@@
for any sequences $r_{n}\to 0,r_{n}>0$ and $h_{n}\to h$ and for $p>0$. Here the derivative $\phi^{\prime}_{\vartheta_{0}}(\cdot)$ does not have to be linear, but is by definition continuous at $h$. If $p=1$ and $\phi^{\prime}_{\vartheta_{0}}(\cdot)$ is linear we say that $\phi$ is Hadamard differentiable at $\vartheta_{0}\in D^{\circ}$ in direction $h\in D$. In Averbukh and Smolyanov, (1968, Theorem 3.2) it is shown that for any compact set $K\subset D$, the Hadamard differentiability at $\vartheta_{0}\in D^{\circ}$ in direction $h$ for all $h\in K$ is equivalent to the compactly differentiability at $\vartheta_{0}$, i.e. there exists a continuous, linear mapping $\phi^{\prime}_{\vartheta_{0}}:D\to E$ with
@@
Similarly, we say $\phi$ is $p$-th order compactly directionally differentiable at $\vartheta_{0}$, if there exists a mapping $\phi^{\prime}_{\vartheta_{0}}:D\to E$ such that
@@
Proposition 4 (Shapiro,, 1990, Proposition 3.3).
Let $p>0$. The $p$-th order Hadamard directional differentiability of $\phi$ at $\vartheta_{0}$ in direction $h$ for all $h\in D$ implies $p$-th order compact directional differentiability of $\phi$ at $\vartheta_{0}$. Conversely, if $\phi$ is $p$-th order compactly directionally differentiability with derivative $\phi^{\prime}_{\vartheta_{0}}$, and if $\phi^{\prime}_{\vartheta_{0}}$ is sequentially continuous, then $\phi$ is $p$-th order Hadamard directionally differentiable at $\vartheta_{0}$.
@@
Another concept, which is stronger than Hadamard differentiability is Fréchet differentiability. We say $\phi$ is Fréchet differentiable at $\vartheta_{0}\in D^{\circ}$, if if there exists a continuous and linear mapping $\phi^{\prime}_{\vartheta_{0}}:D\to E$ such that
@@
Connection to condition (iii) in Proposition 3. In our setting we consider $D=\Theta\times\Theta$ for some Banach space $\Theta$ and are only interested in differentiability at the point $a_{0}=(\vartheta_{0},\vartheta_{0})$ for $\vartheta_{0}\in\Theta^{\circ}$. Consider a loss function $\ell:\Theta\times\Theta\to[0,\infty)$ and set $\phi(a)=\ell(a_{1},a_{2})$ for $a=(a_{1},a_{2})\in\Theta\times\Theta$. Condition (iii) from Proposition 3 requires that for $\Theta\subseteq\mathbb{R}^{d}$ that there exists $\phi^{\prime}_{\vartheta_{0}}:D\to[0,\infty)$ such that for all $h\in\mathbb{R}^{d}$ and compact set $K\subseteq\mathbb{R}^{d}$
@@
$\zeta(t,h)>0,\text{if }t\neq h,\quad\zeta(\lambda t,\lambda h)=\lambda^{p}\zeta(t,h)\text{ and }\zeta(t+a,h+a)=\zeta(t,h),\quad\forall a,t,h\in\mathbb{R}^{d},\lambda>0$
@@
For $d=2$ and $\vartheta_{0}=(0,0)$ the mapping $t\mapsto Z(t)$ is visualized for fixed $Y=(1,1)$. Here $Z$ measures the expected curvature of $\ell$ at $\vartheta_{0}$.
Figure 4: Plot of $(t_{1},t_{2})\mapsto Z(t_{1},t_{2})$ for $Y=(1,1)$. The green arrow is the vector $(1,1)$.
A.2 $L^{p}$-derivative
@@
Let $\mathcal{M}=\{f_{\vartheta}\in L^{p}:\vartheta\in\Theta\}$, with $\Theta\subseteq\mathbb{R}^{d}$, be a parametric family of functions. We say the family $\mathcal{M}$ is $L^{p}$-differentiable at $\vartheta_{0}\in\Theta^{\circ}$, if there exists $\dot{f}_{\vartheta_{0}}:\mathcal{X}\to\mathbb{R}^{d}$ with $\left\lVert\dot{f}_{\vartheta_{0}}\right\rVert\in L^{p}$ and
@@
as $\vartheta\to\vartheta_{0}$. In this form we can see directly that this is the Fréchet differentiability in the space $L^{p}$. We consider three examples of $L^{p}$-differentiability: $L^{2}$-differentiability of the square root of a density, $L^{1}$-differentiability of a density and $L^{1}$-differentiability of a distribution function.
Differentiability in quadratic mean. We say that a parametric family of probability distributions $\{P_{\vartheta}:\vartheta\in\Theta\}$, dominated by some measure $\mu$ on $\mathcal{X}$, is differentiable in quadratic mean at $\vartheta_{0}\in\Theta^{\circ}$, if there is a vector valued function $\ell^{\prime}_{\vartheta_{0}}:\mathcal{X}\to\mathbb{R^{d}}$ such that
@@
Note that the differentiability in quadratic mean is the same concept as the $L^{2}$-differentiability of $\{p^{1/2}_{\vartheta}:\vartheta\in\Theta\}$. Differentiability in quadratic mean implies the existence of the Fisher matrix and
@@
$L^{1}$-differentiability. Differentiability in quadratic mean at $\vartheta_{0}$ implies that the $L^{1}$-differentiability of $p_{\vartheta}$ at $\vartheta_{0}$ and the derivative is given by $\dot{p}_{\vartheta_{0}}=\ell^{\prime}_{\vartheta_{0}}p_{\vartheta_{0}}$, see Liese and Miescke, (2008, Proposition 1.110). So the differentiability in quadratic mean is a stronger property than $L^{1}$-differentiability of the density.
The following proposition gives conditions for $L^{1}$-differentiability of the distribution function $F_{\vartheta}(t)=\int_{-\infty}^{t}p_{\vartheta}(x)dx$. Under the condition that $\mu(\mathcal{X})<\infty$, the result was shown in Proposition B.2 from the supplement of Bernton et al., 2019b . We extend this by allowing $\mathcal{X}$ to be unbounded.
@@
    For any $x\in\mathcal{X}$ the mapping $\vartheta\mapsto p_{\vartheta}(x)$ is continuously differentiable at $\vartheta_{0}\in\Theta^{\circ}$ with derivative $\dot{p}_{\vartheta_{0}}$.
@@
By differentiability in quadratic mean it follows from Liese and Miescke, (2008, Proposition 1.110) that the $L^{1}$-derivative of $p_{\vartheta}$ at $\vartheta_{0}$ is given by $\dot{p}_{\vartheta_{0}}=\ell^{\prime}_{\vartheta_{0}}p_{\vartheta_{0}}$. Define
${\mathcal{F}_{\vartheta}(t)}$ $=\int_{-\infty}^{t}\left\lvert p_{\vartheta}(x)-p_{\vartheta_{0}}(x)-\left\langle\vartheta-\vartheta_{0},\,\dot{p}_{\vartheta_{0}}(x)\right\rangle\right\rvert\mathop{}\!\mathrm{d}\mu(x)$
$=\left\lVert\left(p_{\vartheta}-p_{\vartheta_{0}}-\left\langle\vartheta-\vartheta_{0},\,\dot{p}_{\vartheta_{0}}\right\rangle\right)\mathbb{I}(\cdot\leq t)\right\rVert_{L^{1}}.$
@@
$\left\lVert\vartheta-\vartheta_{0}\right\rVert^{-1}\int_{\mathcal{X}}\left\lvert F_{\vartheta}(t)-F_{\vartheta_{0}}(t)-\left\langle\vartheta-\vartheta_{0},\,\dot{F}_{\vartheta_{0}}(t)\right\rangle\right\rvert\mathop{}\!\mathrm{d}\mu(t)\leq\int_{\mathcal{X}}\frac{{\mathcal{F}_{\vartheta}(t)}}{\left\lVert\vartheta-\vartheta_{0}\right\rVert}\mathop{}\!\mathrm{d}\mu(t).$
@@
$0\leq\lim_{\vartheta\to\vartheta_{0}}\frac{{\mathcal{F}_{\vartheta}(t)}}{\left\lVert\vartheta-\vartheta_{0}\right\rVert}\leq\lim_{\vartheta\to\vartheta_{0}}\frac{\left\lVert p_{\vartheta}-p_{\vartheta_{0}}-\left\langle\vartheta-\vartheta_{0},\,\dot{p}_{\vartheta_{0}}\right\rangle\right\rVert_{L^{1}}}{\left\lVert\vartheta-\vartheta_{0}\right\rVert}=0$
by $L^{1}$-differentiability of $p_{\vartheta}$ at $\vartheta_{0}$. If $\mathcal{X}$ is bounded, we are done. If condition (ii) holds, we use the dominated convergence theorem with the integrable upper bound
$\frac{{\mathcal{F}_{\vartheta}(t)}}{\left\lVert\vartheta-\vartheta_{0}\right\rVert}\leq\sup_{\tilde{\vartheta}\in B_{\delta}(\vartheta_{0})}\left\lVert\left\lVert\dot{p}_{\tilde{\vartheta}}-\dot{p}_{\vartheta_{0}}\right\rVert\right\rVert_{L^{1}}$
@@
As demonstrated in Section 5.2, $L^{1}$-differentiability of the distribution function is closely related to the directional Hadamard differentiability of the 1-Wasserstein distance.